OpenStudy (anonymous):

Find the third iterate, x3, of the function f(x) = 3x + 5 for an initial value of x0 = 1

OpenStudy (anonymous):

Iteration is a process by which the output of one step becomes the input for the next one...so that for f(x) = 2x-5 and x(0) = 3 f(x0) = f(3) = 2(3) - 5 = 1 f(x1) = f(1) = -3 f(x2) = f(-3) = -11 f(x3) = f(-11) = -27

OpenStudy (anonymous):

@aokelley